OpenStudy (anonymous):

Rachel works part-time from home as a data entry professional. She enters data in forms and is paid based on the entry accuracy rate. The table below shows her compensation. tier 1 = $17 tier 2 = $13 tier 3 = $0 Based on her regular performance, the equation below can be used to determine C, the total amount Rachel earns a month entering data, where x represents the total number of forms she completes that month. C = 17(0.9x) + 13(0.05x) + 0(0.05x) What is the minimum number of forms she must complete in a month to earn a total of at least $1,276? 59 70 79 80

OpenStudy (anonymous):

Multiply out the values on the right first and combine xs. \[C = 17(0.9x) + 13 (0.05x) + 0(0.05x) = 15.3x + 0.65x + 0x = 15.95x\] \[1276 = 15.95x\] \[x = 80\]